OpenStudy (anonymous):

use substitution to solve the system of equation 3x-3y=30 -2x+8y=-26 a.4,-6 b.9,-1 c.17, 1 d.infinitely many solutions

OpenStudy (anonymous):

To solve a System of equations by substitution, we find the value of one variable \((\text{say } x)\) in one equation in terms of the other \((y)\). What would the first step be in achieving this?

OpenStudy (anonymous):

im being timed on this test and i really dont know...

OpenStudy (anonymous):

3x-3y=30----- (1) -2x+8y=-26-----(2) _______________ x+5y=4 or,x=4-5y now from equation 1 we can set the value of x 3(4-5y)-3y=30 or,12-15y-3y=30 or,18y=30-12 or,18y=18 or,y=1 now from equation 2 we can find -2x+8=-26 or,-2x=-26-6 or,-2x=-32 or,x=16

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

thank u